# A Path Signature Framework for Detecting Creative Fatigue in Digital Advertising

### v2

@@ 7.1.4 Financial Impact Quantification
The true value of early detection becomes apparent when translated to financial metrics. Using industry-standard CPCs of £0.50-£2.00 (IAB, 2023), a one-day earlier detection for a campaign with 100,000 daily impressions and 1% performance degradation saves:
$\text{Daily Savings}=100,000\times 0.01\times\textsterling 1.25=\textsterling 1,250$
For a portfolio of 100 active creatives, this translates to potential savings exceeding £125,000 daily, justifying the investigation of additional change points flagged by our sensitive detection.
@@ 10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from both simulation and real-world case studies indicate that this signature-based approach provides a more robust, sensitive, and scalable solution than traditional heuristics. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect subtle changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of financial wastage provides a direct bridge from statistical insight to actionable business intelligence.
@@ Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{d}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{d}$.
@@ Appendix C Mathematical Foundations and Theoretical Analysis
C.1 Proof that Signature Distance Captures Meaningful Performance Changes
Theorem 1 (Signature Sensitivity to Path Changes).
@@ C.1 Proof that Signature Distance Captures Meaningful Performance Changes
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|$
where $C>0$ is a constant depending on the path regularity and signature depth $d$.
@@
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. The constant $C$ can be explicitly computed using rough path estimates and depends on the $p$-variation of the paths. ∎
@@ C.2 Theoretical Analysis of Detection Power
Theorem 2 (Power of Signature-Based Change Point Detection).
@@
Under $H_{0}$, the signature distances follow approximately a chi-squared distribution (by the central limit theorem for dependent data). Using concentration inequalities for sub-exponential random variables:
@@
Substituting the lower bound on $\mathbb{E}[D|H_{1}]$ yields the stated result. This shows that detection power increases exponentially with window size and the divergence between pre- and post-change distributions. ∎
@@ C.3 Convergence Properties as Window Size Increases
Theorem 3 (Asymptotic Consistency).
@@
Under regularity conditions (bounded $p$-variation of paths), this process satisfies:
@@
The rate of convergence can be shown to be $O_{p}(w^{-1})$ under additional smoothness assumptions on the underlying process. This demonstrates that larger window sizes provide more accurate change point localisation at the cost of reduced temporal resolution. ∎
@@ C.4 Computational Complexity Analysis
Lemma 1 (Signature Computation Complexity).
For a path of length $w$ in $\mathbb{R}^{d}$ and truncation depth $D$, the computational complexity of computing the truncated signature is:
$\mathcal{O}(w\cdot d^{D}\cdot D)$
However, using the log-signature representation and BCH formula, this can be reduced to:
$\mathcal{O}(w\cdot d^{2}\cdot D^{2})$
@@
The naive computation requires evaluating all iterated integrals up to depth $D$. There are $\sum_{k=1}^{D}d^{k}=\frac{d^{D+1}-d}{d-1}$ components. Each $k$-th level integral requires $O(w)$ operations, giving total complexity $O(w\cdot d^{D})$.
The log-signature approach exploits the group structure of signatures. Using the Campbell-Baker-Hausdorff formula, we can compute log-signatures incrementally with complexity $O(D^{2})$ per increment. For $w$ increments and $d$ dimensions, this yields $O(w\cdot d^{2}\cdot D^{2})$, which is the complexity achieved in our implementation. ∎
@@ C.5 Robustness to Noise and Outliers
Theorem 4 (Noise Robustness).
@@
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C_{d}\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right)$
where $C_{d}$ is a constant depending only on the signature depth $d$.
@@
The sum converges to $C_{d}=e^{\text{Var}(X)}-1$, giving the stated bound. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎

### v3

@@ 7.2.1 Financial Impact Quantification
The true value of earlier detection becomes apparent when translated to operational units. The following calculation is a stylised illustration (not an empirical estimate from our synthetic panel). Given daily impressions $I$, benchmark cost-per-click $CPC_{bench}$, and an absolute CTR shortfall of $\Delta CTR$ relative to a benchmark period, the implied opportunity cost per day is
$\text{Opportunity Cost per day}=I\,\Delta CTR\,CPC_{bench}.$
This mapping is provided only to translate lead-time into operational units; it is not intended as a validated estimate of platform spending dynamics.
@@ 10 Conclusion
In this paper, we have introduced a novel methodology for identifying and quantifying advertising creative fatigue by applying the mathematical framework of path signatures. Our findings from a synthetic panel evaluation indicate that this signature-based approach can provide early warnings under an explicit operational onset definition, while highlighting a practical trade-off between early warning and alert burden. By treating the performance time series as a geometric path and using its signature as a rich feature descriptor, our method can detect changes in the underlying dynamics of creative performance that simpler models may miss. The subsequent quantification of this fatigue in terms of opportunity cost provides a bridge from statistical outputs to operational prioritisation.
@@ Appendix A A Deeper Dive into Rough Path Theory
Rough path theory provides a mathematical framework for describing the interactions between complex, oscillating systems. At its core is the concept of the path signature, which serves as a faithful, non-parametric summary of a path. For a continuous path $X_{t}$ in $\mathbb{R}^{m}$, its signature $S(X)$ is an infinite sequence of tensors, living in the tensor algebra space over $\mathbb{R}^{m}$.
@@ Appendix C Mathematical Foundations and Theoretical Analysis
This section provides informal theoretical support for why signature distances are sensitive to changes in the underlying trajectory. For readability, we present proof sketches rather than fully detailed proofs. We assume paths are continuous and of bounded variation (or, more generally, finite $p$-variation) so that truncated signatures are well-defined. To avoid notational ambiguity, we use $m$ to denote the path dimension and $d$ to denote the signature truncation depth.
C.1 Signature distance sensitivity
Theorem 1 (Signature sensitivity to path changes (proof sketch)).
@@ C.1 Signature distance sensitivity
$\|S(X)_{[t_{1},t_{2}]}-S(Y)_{[t_{1},t_{2}]}\|\geq C\cdot\sup_{t\in[t_{1},t_{2}]}\|X(t)-Y(t)\|,$
where $C>0$ depends on regularity assumptions and the truncation depth $d$.
@@
For paths with different geometric properties (e.g., one declining, one stable), the higher-order signature terms amplify these differences, providing the stated lower bound. A rigorous statement requires explicit norms and assumptions (e.g., bounded $p$-variation) and follows from continuity estimates for truncated signatures. ∎
@@ C.2 Theoretical Analysis of Detection Power
Theorem 2 (Detection power (heuristic bound)).
@@
Under $H_{0}$, the signature distances concentrate around their mean; a complete derivation depends on dependence assumptions in the time series and on the distribution of the signature estimator. Using concentration-style reasoning:
@@
Substituting lower bounds on $\mathbb{E}[D|H_{1}]$ yields the stated qualitative dependence. The intent is to highlight that detection improves with larger window sizes and with greater separation between pre- and post-change regimes. ∎
@@ C.3 Convergence Properties as Window Size Increases
Theorem 3 (Asymptotic consistency (informal)).
@@
Under regularity conditions (e.g., bounded $p$-variation and appropriate mixing), this process satisfies:
@@
The rate of convergence depends on additional assumptions and is stated here for intuition only. Larger window sizes provide more stable estimation at the cost of reduced temporal resolution. ∎
@@ C.4 Computational Complexity Analysis
Lemma 1 (Signature computation complexity (order of growth)).
For a path of length $w$ in $\mathbb{R}^{m}$ and truncation depth $d$, the number of truncated signature terms grows as $\sum_{k=1}^{d}m^{k}$. A naive computation scales approximately as:
$\mathcal{O}(w\cdot m^{d}\cdot d).$
In practice, log-signature representations reduce constant factors substantially; for fixed $m=2$ and small $d$ (as used in our experiments), the overall detection pipeline is effectively linear in the number of time points.
@@
The dominant cost is evaluating iterated integrals (or equivalent algebraic updates) over $w$ increments. Since the number of terms grows exponentially in $d$, practical implementations fix $d$ (e.g., $d\in\{3,4\}$) and treat it as a small constant. ∎
@@ C.5 Robustness to Noise and Outliers
Theorem 4 (Noise robustness (informal)).
@@
$\mathbb{E}[\|S(\tilde{X})-S(X)\|^{2}]\leq C\cdot w\cdot\sigma^{2}_{\epsilon}\cdot\left(1+O(w^{-1/2})\right),$
where $C$ depends on the truncation depth $d$ and regularity of $X$.
@@
The sum converges to a finite constant under bounded-variation assumptions, giving the stated qualitative dependence. This shows that the signature-based method has linear sensitivity to noise variance, with the normalisation step providing additional robustness. ∎